Although appealing to more refined tastes, art as a collectible has not always performed so profitably During 2003, an auction house sold a sculpture at auction for a price of $10,281,500. Unfortunately for the previous owner, he had purchased it in 1998 at a price of $12,317,500. What was his annual rate of return on this sculpture? (Negative amount should be indicated by a minus sign. Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) Annual rate of return 0
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

Hence

10,281,500=12,317,500(1+r/100)^5

(10,281,500/12,317,500)^(1/5)=(1+r/100)

(1+r/100)=0.9645

r=(0.9645-1)*100

=(3.55%)(Negative).